# Programmable photonic waveguide arrays: opportunities and challenges

### v1

@@ III.1 Overview of continuously coupled waveguide systems
Programmable waveguide arrays are assembled by adding precisely positioned microtuners to the static WA structures. These microtuners placed around the waveguides allow modification of the device’s Hamiltonian by applying control pulses. The design of microtuners can vary depending on the selected tuning mechanism and material platform, which is discussed in the following content. An ideal PWA enables precise control over each term in its Hamiltonian, thereby rendering it fully programmable. Consequently, the unitary transformation implemented by such devices can be controlled. A conceptual schematic adapted from the recently demonstrated PWA device [42] is shown in Fig. 3.
III.2 Theoretical model
@@ III.3 Material platforms
III.4 Circuit layout and waveguide geometry optimization
@@ IV Characterization and control
IV.1 Device characterization
The process of device characterization involves measuring the device to establish a mapping between its control parameters and the corresponding Hamiltonian or unitary transformation matrix [186]. Characterizing a PWA device to explore its achievable unitary space involves modeling the device and fitting the model using numerous measurements. These measurements can be performed using either classical light [187, 188, 189] or quantum light via HOM measurements [190, 191, 192], similar to the process used in mesh scheme characterizations. However, the mathematical approach to constructing the device model differs, as PWAs do not allow completely decoupled waveguides. Therefore, machine learning-based methods are commonly employed [188, 112], typically relying on a predefined physical model, such as the one discussed in Sec. III.2. Other approaches, such as optimization techniques [193, 194] or supervised learning algorithms [195], can also be adapted to characterize PWAs.
@@ IV.2 Device control
A study with a 3-PWA system [112] revealed that the method in Sec III.2 is incomplete for modeling a fabricated physical device. This work shows that by relaxing the tri-diagonal Hamiltonian assumption and allowing for a complex-valued Hamiltonian, the fitting performance of the model is improved significantly. This is due to the fact that the fabricated device is non-homogeneous along the propagation direction due to either material non-homogeneity or imperfect electrode alignment. This suggests that a fabricated device deviates notably from the model of an ideal device.
@@
Figure 4: Cross-section of a two-dimensional PWA concept design. The optical modes are arranged in a hexagonal pattern, as indicated by dashed lines. Four tuners between neighboring optical modes control their coupling coefficients, while six tuners placed around an optical mode adjust their propagation coefficients. The substrate is made of an ideal isotropic material, such as one with a homogeneous high electro-optic coefficient in all directions.
@@ V Conclusion and outlook
Programmable waveguide arrays integrate the unique properties of static waveguide arrays, including “always-on” Hamiltonian implementation, compactness, and robustness. The reconfigurability of PWAs, which allows direct control over device Hamiltonians parameters, enables the realization of diverse functions within a single platform, ranging from light propagation experiments and quantum walks to topological photonics. This versatility removes the need to fabricate multiple, application-specific static devices. PWA-based architectures enable the implementation of arbitrary unitary transformations as well as the time evolution of time-dependent Hamiltonians. This capability extends their functionality beyond universal multiport interferometers, unlocking a wide range of applications in communication, sensing [204], and hybrid digital-analog computation. Moreover, robust protocols harnessing the advantages of PWAs in implementing topological effects are likely to be developed for advanced fields such as quantum computing and neuromorphic computing.
The theoretical analysis presented in paper [43] was based on the constrained tridiagonal Hamiltonian (Eq. 1) under a first-order approximation. For physical materials, the study also reveals that the performance of a single-section PWA is inherently constrained by the tridiagonal approximation, allowing only the approximation of tridiagonal unitaries. Future work could focus on deriving tighter bounds that align more closely with numerical optimization results. Additionally, incorporating higher-order coupling terms beyond nearest-neighbor interactions [150] could reduce the number of sections needed to achieve a desired fidelity level.
Prioritizing the design and fabrication of PWAs on highly tunable material platforms could enable the development of more compact devices, especially for future upscaling. Compared to MZI-based structures, the circuit size of PWA-based processors could be significantly reduced by eliminating all bends, though further work is needed to confirm this. In this context, PWA-based architectures hold the potential to sustain the growth trend illustrated in Fig. 1.
Furthermore, the programmability of a two-dimensional waveguide array structure could be achieved by embedding microtuner units alongside the waveguides [205], as illustrated in the conceptual design shown in Fig. 4. This design allows for arbitrary tuning of the device’s Hamiltonian parameters and enables coupling between more than two adjacent waveguides, in contrast to the planar design shown in Fig. 3. As a result, this approach has the potential to create a more compact device enabling the possible implementation of arbitrary unitaries with less number of sections compared to the equivalent planar structure. However, tuner placement remains a significant challenge when designing two-dimensional PWAs and is heavily constrained by the material platform.
Future work could also be extended to implementing Haar random unitaries with PWAs, which is the foundation of complexity conjectures [206, 207]. As the resources required for the characterization and control of PWA systems increase with their dimension, developing more sophisticated modeling and characterization methods becomes essential to fully harness their potential.

### v3

@@ II.1 Overview of continuously coupled waveguide systems
Programmable waveguide arrays are assembled by adding precisely positioned microtuners to the static WA structures. These microtuners placed around the waveguides allow modification of the device’s Hamiltonian by applying control pulses. The design of microtuners can vary depending on the selected tuning mechanism and material platform, which is discussed in the following content. An ideal PWA enables precise control over each term in its Hamiltonian, thereby rendering it fully programmable. Consequently, the unitary transformation implemented by such devices can be controlled. A conceptual schematic of a PWA, adapted from the recently demonstrated device [40] is shown in Fig. 2.
II.2 Theoretical model
@@ II.3 Material platforms
II.4 Circuit layout and waveguide geometry optimization
@@ III Characterization and control
III.1 Device characterization
The process of device characterization involves measuring the device to establish a mapping between its control parameters and the corresponding Hamiltonian or unitary transformation matrix [102]. Characterizing a PWA device to explore its achievable unitary space involves modeling the device and fitting the model using numerous measurements. These measurements can be performed using either classical light [103, 104, 105] or quantum light via HOM measurements [106, 107, 108], similar to the process used in mesh scheme characterizations. However, the mathematical approach to constructing the device model differs, as PWAs do not allow completely decoupled waveguides. Therefore, machine learning-based methods are commonly employed [104, 59], typically relying on a predefined physical model, such as the one discussed in Sec. II.2. Other approaches, such as optimization techniques [109, 110] or supervised learning algorithms [111], can also be adapted to characterize PWAs.
@@ III.2 Device control
A study with a 3-PWA system [59] revealed that the method in Sec II.2 is incomplete for modeling a fabricated physical device. This work shows that by relaxing the tri-diagonal Hamiltonian assumption and allowing for a complex-valued Hamiltonian, the fitting performance of the model is improved significantly. This is due to the fact that the fabricated device is non-homogeneous along the propagation direction due to either material non-homogeneity or imperfect electrode alignment. This suggests that a fabricated device deviates notably from the model of an ideal device.
@@
IV Translating potential applications to the programmable paradigm
The following sections examine key domains where PWAs could significantly expand functionality, efficiency, or usability, even if full demonstrations are yet to be realized. Each application area is introduced through its static WA heritage or MZI-based implementation, followed by an exploration of how programmability could enable new capabilities or resolve existing bottlenecks.
IV.1 Universal multiport interferometers
Application Overview:
Universal multiport interferometers enable arbitrary linear unitary matrix transformations between multiple optical channels. They have extensive applications in both classical and quantum information processing [17, 36], which will be discussed in more detail in the following content. In addition to MZI-based mesh structures, which are considered discrete or digitized, there has been growing interest in leveraging continuously coupled waveguides or multimode interferometers (MMIs) as error-robust and more scalable alternatives to realize universal multiport interferometers [120, 121]. In 2017, a hybrid structure based on cascading MMIs and phase shifters was proposed to implement unitary transformations [122]. This structure was further investigated in [123], where several transformations were numerically studied. Experimental demonstrations of such an architecture have also been performed [124, 125]. Enhanced scalability and error tolerance have also been demonstrated for large-scale circuits [126]. The realization of multiport transformations using structures with partially programmable waveguide arrays has also been proposed [127].
Potential of PWAs:
Structures based on cascaded PWAs take full advantage of continuously coupled waveguides in integrated photonics. They have also been recently rigorously proven to be universal [44], addressing a long-standing question raised in previous studies. The work introduced a decomposition algorithm for realizing any unitary matrix using PWAs with guaranteed error bounds. For $U(2)$, a maximum of 4 sections are needed, which is the exact decomposition. The required number of sections $n$ scales as $O(N^{3}k)$, with the error scaling as $O(N^{3}/k)$, where $k$ is the Trotterization number. While this result suggests a high number of sections may be required to achieve low fidelities, the study demonstrates that, in practice, the number of sections needed is significantly smaller than the theoretical bound. By leveraging realistic materials parameters and numerical optimization to compile the unitary, these requirements become practically achievable. Therefore, cascaded PWAs are functionally equivalent to the widely adopted MZI-mesh structures [18, 19], while they eliminate the need for bending. This unique feature can potentially lead to significant reductions in optical losses and circuit footprint.
IV.2 Quantum computing
Application Overview:
Quantum computers promise to solve computational problems more efficiently than classical systems [128]. In 1982, Feynman proposed using controllable quantum systems to simulate the dynamics of other systems [129], now known as analog quantum computing [130, 131]. This approach allows the simulator’s Hamiltonian to be controlled to match the target system, offering greater error robustness, less demanding hardware, and better scalability compared to universal digital architectures [132, 133, 134]. Waveguide arrays, which naturally implement time evolution governed by their Hamiltonian, have been used to simulate optical analogs of solid-state systems [135]. By varying waveguide coupling along the propagation length and integrating control electrodes to tune coupling strength, PWAs can simulate time-dependent Hamiltonians for analog quantum simulations. Integrated photonics is a promising platform for gate-based quantum circuits [136, 99]. Digital quantum computing [137], where qubits evolve through single- and two-qubit gates, is universal [138] and can be implemented in photonic systems using mesh structures of $2\times 2$ beam splitters and phase shifters [18, 19, 20]. While static, continuously coupled waveguides have been used to implement quantum gates [101, 38], digital quantum circuits require reconfigurable systems.
Potential of PWAs:
Recently, controllable single-qubit gates have been demonstrated [41], and cascaded PWAs offer a pathway to universal digital quantum circuits [44]. A PWA-based digital quantum simulator could implement Trotterization steps by discretizing time evolution into small unitary operations [138]. By tuning the waveguide coupling and interaction parameters, this setup can approximate the continuous evolution of a system’s Hamiltonian, simulating digital quantum gates effectively. This approach provides a scalable route to implement digital quantum algorithms using cascaded waveguide arrays. By supporting both analog and digital quantum information processing, cascaded PWAs offer a unique capability in photonics. Digital-analog quantum computing (DAQC), which combines the strengths of both approaches, has been proposed to address the challenges of the noisy intermediate-scale quantum era [139]. Using analog methods for continuous evolution and digital precision for error correction, DAQC enhances efficiency, error tolerance, and scalability, offering a promising path to solve complex quantum problems with fewer resources.
IV.3 Photonic neural networks
Application Overview:
Potentials of PWAs:
Programmable waveguide arrays present a compelling alternative hardware platform for implementing photonic neural networks. Compared to MZI-based or hybrid MMI–phase shifter architectures, PWAs may more easily scale to larger input dimensions, support multi-wavelength operations, and reduce overall optical losses. Furthermore, due to the universality and robustness of PWAs, the training of PWA-based PNNs can be streamlined with a simplified calibration process. Additionally, the continuously coupled waveguides in PWAs provide unique interfaces for integrating nonlinear activation elements.
IV.4 Topological photonics
Application Overview:
Topological photonics has become a vibrant and rapidly evolving field over the past decade, drawing inspiration from the discovery of quantum Hall effects and topological insulators in condensed matter physics [14, 42, 147]. WAs, with their ability to emulate tight-binding Hamiltonians, have proven to be powerful simulators of electron behavior in such systems [148]. This has enabled the experimental exploration of a wide range of topological phenomena, including the first demonstration of parity-time (PT) symmetry in optics [149] and photonic Floquet topological insulators [58]. These topological systems usually exhibit intrinsic robustness to disorder [150], making them promising for practical applications in areas such as optical isolation [151, 152], communications [153], spintronics [154], and even quantum computing. A wide range of topologically protected quantum states has been realized in WA-based platforms [155, 156, 39, 37, 157, 158, 159, 160, 161, 162], affirming the potential of these systems for fundamental science and future technologies.
Potentials of PWAs:
PWAs are poised to significantly expand the capabilities of topological photonic systems. Unlike static WAs, PWAs allow the exploration of multiple topological phases within a single device. This programmability also opens the door to dynamically probing robustness by introducing controlled disorder or tunable defects [40]. Additionally, with reconfigurable coupling and phase shifts, PWAs could provide a flexible testbed for implementing topological optical isolators, as well as exploring emergent phenomena in disordered or time-varying systems [163, 164].
Furthermore, the modularity of PWAs suggests a potential route toward cascading multiple two-dimensional arrays with integrated dynamic modulation. Such architectures may serve as compact alternatives to helically curved designs, enabling photonic simulations of time-dependent Hamiltonians while mitigating bending-induced losses [165, 166]. With these features, PWAs could be well-positioned to serve as a flexible and scalable platform for next-generation topological photonic devices.
IV.5 Nonlinear phenomena
Application Overview:
Nonlinear optical effects have significantly expanded the functional landscape of integrated photonics, particularly in coupled waveguide systems. Early work began with nonlinear directional couplers [167] and soon extended into WAs [168], where nonlinear interactions gave rise to spatial solitons [57, 24], nonlinear Bloch oscillations [28, 29], and complex dynamics in two-dimensional lattices [25, 27]. These studies have contributed to advances in pulse shaping [169], all-optical switching [170]
In recent years, the integration of nonlinear optics with topological photonics has opened even more compelling avenues. Nonlinear WAs have enabled the realization of self-localized topological edge states [171, 172], topological insulator lasers that maintain robustness under fabrication disorder [173], and chip-scale generation of topologically protected entangled photon pairs via SPDC and four-wave mixing [174, 175, 176, 177]. These breakthroughs highlight the growing potential of nonlinear photonics as a foundation for both classical and quantum optical technologies.
Potentials of PWAs:
PWAs offer a promising path to dynamically enhance and control nonlinear phenomena in integrated photonics. Unlike static WAs, PWAs enable real-time control over coupling and phase parameters, making it possible to dynamically shape nonlinear interactions. This tunability provides a new level of flexibility for controlling soliton dynamics, optimizing nonlinear Bloch oscillations, and adaptively shaping pulses for ultrafast optics or optical computing. Moreover, when built from or integrated with materials exhibiting strong second- or third-order nonlinearities, PWAs could become tunable platforms for efficient wavelength conversion, light filtering, and dynamic routing in nonlinear circuits.
The ability to combine nonlinearity with topological protection in a programmable platform is particularly powerful. With PWAs, one could study the emergence and control of nonlinear topological edge states under varying interaction strengths, or explore regimes where topological protection competes or cooperates with nonlinear localization. Reconfigurability also enables new approaches to simulate nonlinear-driven phase transitions and dynamically explore higher-order topological phenomena that are otherwise difficult to access with static systems.
IV.6 Other potential emerging applications
Additional potential applications of precisely engineered device Hamiltonians and lengths include quantum walks [178, 35, 179], simulating non-Markovian giant atom decay [180] and higher-dimensional systems [181], emulating and harnessing non-Abelian braiding  [182, 183], generating W states [184, 185], and achieving perfect state transfer [186]. The reconfigurability of PWAs offers a unified platform for realizing these applications with enhanced precision, scalability, and adaptability. Time-dependent Hamiltonians could potentially be realized by cascading PWAs, with each section implementing a different Hamiltonian [58, 166, 165]. Moreover, the incorporation of gain and loss in PWAs can be achieved by applying various control mechanisms to waveguide tuners, including the creation of defect waveguides [187, 188]. This setting enables the realization and control of non-Hermitian dynamics [189], unlocking phenomena such as parity-time symmetry [149] and topological phase transitions [190], as well as applications like the generation of topologically protected bound states [191].
@@ IV.6 Other potential emerging applications
Figure 4: Cross-section of a two-dimensional PWA concept design. The optical modes are arranged in a hexagonal pattern [58], as indicated by dashed lines. Four tuners between neighboring optical modes control their coupling coefficients, while six tuners placed around an optical mode adjust their propagation coefficients. The substrate is made of an ideal isotropic material, such as one with a homogeneous high electro-optic coefficient in all directions.
@@ V Conclusion and outlook
Programmable waveguide arrays combine the Hamiltonian richness and compact form factor of static waveguide arrays with the versatility of reconfigurable photonic platforms. Their ability to implement “always-on” Hamiltonians with tunable parameters enables a wide range of applications, including quantum walks, analog simulations, topological photonics, and photonic neural networks. Unlike mesh-based programmable architectures, PWAs eliminate the need for bending and beam splitting, offering potential advantages in optical loss, circuit footprint, and fabrication tolerance.
Recent demonstrations have shown that cascaded PWAs are capable of implementing arbitrary unitary transformations, including time-dependent Hamiltonian evolution, providing a flexible framework for quantum information processing, reconfigurable signal processing, and hybrid digital-analog computation. This versatility may unlock high-dimensional encoding schemes in communications [192], robust sensing with programmable dynamics [193, 194], and new protocols for quantum and neuromorphic computing.
The theoretical framework presented in [44] establishes the universality of PWAs under a constrained tridiagonal Hamiltonian model (Eq. 1), though practical performance is inherently limited by fabrication variability and first-order approximations. Incorporating higher-order coupling terms [64] or transitioning to multi-dimensional geometries may reduce the number of sections required to achieve target fidelities. Two-dimensional PWA structures could be enabled by integrating liquid metals as tunable elements [195], as illustrated in Fig. 4, offering a pathway toward more compact architectures with richer connectivity and enhanced control.
To advance towards scalable and deployable PWA-based systems, the field must now establish clear quantitative performance benchmarks. We identify four immediate priorities:
  1. 1.
    Increasing the number of precisely programmable waveguides and sections beyond current prototypes, typically limited to 3–11 waveguides and a single section, remains a major design and fabrication challenge.
  2. 2.
    Defining the achievable level of control per waveguide and per section, including independent tuning of coupling and phase parameters, is essential for establishing the functional limits of PWA architectures.
  3. 3.
    Systematically reporting key performance indicators–such as insertion loss, control voltage range, and coupling tunability–is needed to enable meaningful cross-platform comparisons (e.g., with MZI-based meshes).
  4. 4.
    Evaluating the engineering gap between academic demonstrations and deployable systems across target applications will help identify the dominant bottlenecks and guide technology maturation strategies.
These benchmarks will be essential for advancing the technology, comparing material platforms, and determining whether PWAs are ready to support real-world applications or remain limited to proof-of-concept experiments.
Future work may also explore the implementation of Haar-random unitaries in PWAs, which form the basis of complexity conjectures in photonic quantum computing [196, 197]. As the number of tunable parameters grows, the need for scalable characterization and control techniques becomes increasingly urgent. Advances in data-driven modeling, in-situ feedback, and high-density tuning elements will be instrumental in overcoming these challenges.
Overall, programmable waveguide arrays offer a promising new paradigm for reconfigurable integrated photonics. As control fidelity, circuit complexity, and integration improve, PWAs are poised to play a central role in scalable quantum technologies, adaptive optics, and photonic information processing.